OpenStudy (anonymous):

Divide. 9.43 divided by 10^-5 94.3 94300 943000 0.0000943

OpenStudy (anonymous):

@Mathbreaker @alakara_kibo

OpenStudy (anonymous):

D

OpenStudy (anonymous):

9.43/10^-5=9.43*10^5=?

OpenStudy (anonymous):

The answer is not D actually, because you are dividing with 1/10^5 technically speaking. Therefore, its actually 9.43/1/10^5 which can also be written as 9.43*10^5, that gives you 943000. correct me if im wrong please but i think thats what it is. :)